# Probabilistic Modeling of Antibody Kinetics Post Infection and Vaccination: A Markov Chain Approach

### v1

@@ 7 Example applied to population transmission
While the previous section was focused on antibody responses in individuals, the transmission of a disease through a population is another important factor that governs the total antibody distribution in a population over time. We now turn our attention to the different possible personal event sequences for individuals as an infection and its vaccination transmits through a population. We consider three separate time-homogeneous transition matrices and analyze their behavior over $50$ time steps, or around $2$ years.
We first consider $S_{1}$ representing a disease spread with a rate of infection that is a magnitude higher than that of vaccination. Recall that our states are ordered as $N,\mathcal{I},\mathcal{V},\mathcal{I}^{\prime},\mathcal{V}^{\prime}$ in all transition probability matrices with the column representing the transition out of the corresponding state.
[…]
Figure 9: Histograms of counts of total infections, vaccinations, previously infected status, previously vaccinated status, and naïve status for a population with $1000$ individuals with a time-homogeneous transition matrix $S_{1}$, per $50$ time steps, i.e., approximately two year period.
@@
We observe (Fig. 9) that due to larger probabilities of infection, the median number of infections per individual is $4$ and the median previously infected status is maintained for $35$ time steps. In contrast, the vaccinations have medians of 00 and naïve status is maintained for a median of $8$ time steps. Some representative individual event sequences are demonstrated in Fig. 10, each row represents a personal trajectory for an individual, the colors represent the immune status for the individual at the time step corresponding to the column.
@@
Figure 10: Three representative sequences for a population with $1000$ individuals with a time-homogeneous transition matrix $S_{1}$, for $50$ time steps, i.e., approximately two year period. Here, light yellow represents the naïve state (1, $N$), red and pink represent the newly and previously infected states (2, $\mathcal{I}$ and 4, $\mathcal{I}^{\prime}$), and blue and light blue represent the newly and previously vaccinated states (3, $\mathcal{V}$ and 5, $\mathcal{V}^{\prime}$) respectively.
We increase the probability of transition to naïve from both previously infected and previously vaccinated status to obtain $S_{2}$.
[…]
@@
This emulates diseases with shorter-term immunity and thus a quicker return to the naïve status; see Fig. 11. As anticipated, this increases the likelihood of sequences with higher number of immune events; see Fig. 12.
Figure 11: Histograms of counts of total infections, vaccinations, previously infected status, previously vaccinated status, and naïve status for a population with $1000$ individuals with a time-homogeneous transition matrix $S_{2}$, per $50$ time steps, i.e., approximately two year period.
@@
Figure 12: Five representative sequences for a population with $1000$ individuals with a time-homogeneous transition matrix $S_{2}$, for $50$ time steps, i.e., approximately two year period. Here, light yellow represents the naïve state (1, $N$), red and pink represent the newly and previously infected states (2, $\mathcal{I}$ and 4, $\mathcal{I}^{\prime}$), and blue and light blue represent the newly and previously vaccinated states (3, $\mathcal{V}$ and 5, $\mathcal{V}^{\prime}$) respectively.
[…]
As expected, the number of vaccinations per individual and stay in previously vaccinated states rises for $S_{3}$ as compared to $S_{2}$ (see Fig. 13). Moreover, the time spent in naïve or naïve-like states decreases. We see that individuals frequently move from a previously infected/vaccinated state to another immune event, (see Fig. 14.)
Figure 13: Histograms of counts of total infections, vaccinations, previously infected status, previously vaccinated status, and naïve status for a population with $1000$ individuals with a time-homogeneous transition matrix $S_{3}$, per $50$ time steps, i.e., approximately two year period.
Figure 14: Five representative sequences for a population with $1000$ individuals with a time-homogeneous transition matrix $S_{3}$, for $50$ time steps, i.e., approximately two year period. Here, light yellow represents the naïve state (1, $N$), red and pink represent the newly and previously infected states (2, $\mathcal{I}$ and 4, $\mathcal{I}^{\prime}$), and blue and light blue represent the newly and previously vaccinated states (3, $\mathcal{V}$ and 5, $\mathcal{V}^{\prime}$) respectively.
@@ 8 Discussion
In this section, we discuss a generalization of the two-event time-inhomogeneous model, more on the general time-inhomogeneous model, and then investigate interesting questions related to antibody data: namely, when to measure the event of vaccination for multi-dose series, and the variability of antibody response across a population. We include a comparison of the two datasets used in our validation of our methods, and finish with limitations and extensions of our work, as well as implications for immunologists.
@@ 8.2 General time-inhomogeneous model
Towards constructing the conditional probability models in the general setting of the time-inhomogeneous model, we discuss an important check. We want to ensure that the number of high-probability sequences with a large number of events is vanishingly small to reflect real world situations, as a person will not realistically become reinfected or revaccinated at every point in time. Notice that with the possibility of either infection or vaccination at each point in time, we obtain $2^{T}$ possible event sequences over $T$ time steps. In reality however, we can constrain this number by quite a bit. We note that as class $\mathcal{I}$ necessarily transitions to class $\mathcal{I}^{\prime}$ and class $\mathcal{V}$ to class $\mathcal{V}^{\prime}$, there are at the most $\lfloor T/2\rfloor$ events possible before time $T$, which means there are approximately $2^{\lfloor T/2\rfloor}$ ways for events to occur before time $T$. We recall also that time will likely be batched by several weeks, so $T=10$ might mean 210 days if $\Delta T=21$ days or 3 weeks, which is a reasonable time spacing. Then for this example, there would be $322^{5}=$ 32 possible sequences in 210 days, which is a manageable number. We can further constrain this set by calculating the probability of obtaining such a sequence as a product of the associated transition probabilities. What we need to additionally consider are the days spent between two immunological events in the previously infected or previously vaccinated state as this affects the antibody response distribution. As an illustration, consider $T=6$. For simplicity, let us assume that $z_{i}\in\{\mathcal{I},\mathcal{V}\}$. The 43 total trajectories are:
@@
8.3 When to measure the event of vaccination?
Because some vaccines are given as a multi-dose sequence, and an individual is not considered fully vaccinated until the sequence is finished, it is interesting to study the effect on our models of setting $t=0$ to represent the day of the first dose of vaccination, rather than the second, as we have done in Section 6. The relative days between the first and second dose of vaccination were available for Dataset 1. We redid our models using dose 1 as $t=0$ (call this Vax Model Dose 1); the results are shown in Table 3 and Fig. 15. We note that this allows for more data to be used; in Section 6, only measurements post dose 2 are relevant for the model (call this Vax Model Dose 2; see Fig. 3).
Compared to Vax Model Dose 2, $\theta_{\mathcal{V}_{1}}$ and $\phi_{\mathcal{V}_{1}}$ are an order of magnitude smaller for Vax Model Dose 1; as a result, the peak and decay of Vax Model Dose 1 are smaller and start later, around 60 days as opposed to 10. Ultimately, we are observing the same peak in the two models, which is the response to the second immunization. More of the increase in antibody response is visible when including data after only the first dose in the model; this is paralleled by the infection model (see Fig. 3). Perhaps reassuringly, this modeling choice for the first vaccination event has minimal effect on the model for the booster dose; the parameter values are relatively similar for Vax Model Dose 1 and Vax Model Dose 2.
Figure 15: Log-transformed antibody measurements from the vaccinated (V) and boosted (VV) populations from Dataset 1 with corresponding probability models.
Vaccination 44.1 2.17 2.56
Vaccination,
booster
44.1 2.17 2.56 155 12.6 1.51
Table 3: Optimal parameters from maximum likelihood estimation for Dataset 1 for measurements corresponding to vaccination as the first event. Grayed out numbers indicate variables that are part of the model but are not part of the optimization for that data. Time is measured since first vaccination dose.
@@ 8.4 Personal trajectories of interest
Figure 16: Log-transformed antibody measurements from the infected (I) (a) and vaccinated (V) (b) populations from Dataset 1 with personal trajectories of interest with corresponding probability models.
We study personal trajectories of interest for a few subjects from both datasets. Although the models shown in Section 6 take into account the time between an event and the day of measurement, it would be difficult to visualize this time span in the two-event figures, as there is often no associated measurement value for the day of the event and the extra information may crowd already densely packed data points. When studying personal trajectories, one option is to mark the line connecting available measurements with an additional marker symbolizing the day of the event. Here, while studying personal trajectories one by one, we mark the date of vaccination (after infection or as a booster dose) as a vertical, dashed line. To discuss individual subjects, we standardize and anonymize the IDs by dataset (D1 or D2 for Dataset 1 or Dataset 2) and appearance number in the following figures by dataset (e.g., D1-1).
In Fig. 16, we plot the single-event models for infection and vaccination with representative and outlier personal trajectories of interest for Dataset 1. For infection, in Fig. 16a, we see that Subject D1-5 maintains a high response and Subjects D1-1, D1-3, and D1-4 illustrate medium responses after the initial acute phase that correspond to relatively high probability values for all involved measurements. However, Subject D1-1 shows a rapid but medium response initial increase, Subject D1-4 has a rapid, large initial increase, and Subject D1-3 has a slower and medium response initial increase in antibody level. Subject D1-2 has a rapid, high initial response and then a quick decay to a below average antibody level.
For vaccination, in Fig. 16b, we see that Subject D1-8 shows a medium response after the initial vaccination. Subjects D1-10 and D1-11 are low responders, as Subject D1-11 never mounts much of an antibody response at all, and Subject D1-10’s response falls below the lowest model contour before 200 days after vaccination. Subjects D1-7 and D1-9 have relatively high initial responses, but Subject D1-7’s antibody level falls much more quickly than Subject D1-9. Subject D1-6 exhibits a medium initial response, but this is sustained better than many subjects beyond 100 days, and then we speculate that they experienced a missed infection sometime before their next measurement (over 400 days after vaccination).
Figure 17: Log-transformed antibody measurements from the infected then vaccinated (VI) population from Dataset 1 with personal trajectories of interest with corresponding probability models. The relative date of vaccination is marked with a vertical, white dashed line.
In Fig. 17 we plot individualized two-event models for vaccination after infection for three subjects of interest. Subject D1-12 has a representative relative time between infection and vaccination (182 days), and exhibits a standard increase in antibody response after vaccination. Both measurement values fall on the higher side of observed values. Subject D1-13 exhibits a strong initial response to infection and a typical decay. Nearly 400 days elapse between their infection and vaccination; likely as a result, they show a strong and sustained response to vaccination. Subject D1-14 is an outlier for their infection response, displaying a very low antibody level almost 200 days after infection, but perhaps in response, they exhibit a very strong increase in antibody level after vaccination.
Figure 18: Log-transformed antibody measurements from the vaccinated then boosted (VV) population from Dataset 1 with personal trajectories of interest with corresponding probability models. The relative date of the booster dose is marked with a vertical, white dashed line. See Supplementary Material for accompanying video.
In Fig. 18 we plot individualized two-event models for a booster dose after vaccination for six subjects of interest. Subject D1-15 exhibits a medium response to both infection and vaccination, with representative decay patterns that are well captured by our model. This is true for Subject D1-16 for the measurements up until the booster dose, roughly 100 days after which we speculate that a breakthrough infection occurred, since the subject’s antibody response spikes again. We observe that Subject D1-17 mounted a relatively low response to the first vaccine sequence, but a stronger response was induced by the booster dose. Subject D1-18 exhibits a low initial vaccine response that drops quickly, followed by a medium response to the booster that also rapidly drops. Finally, Subjects D1-20 and D1-19 show low and high antibody responses to both the initial and booster vaccinations, respectively, with a particularly high and sustained response to the booster dose for Subject D1-19.
@@
Figure 19: Log-transformed antibody measurements from the infected (I) and vaccinated (V) populations from Dataset 2 with personal trajectories of interest with corresponding probability models. Measurements with recorded monotonically decreasing OD values have been AUC-extrapolated using the next titration level.
@@
In Fig. 19 we plot representative and outlier single-event personal trajectories with the corresponding probability models for Dataset 2. In Fig. 19(a), Subject D2-5 exhibits a high response, Subject D2-3 demonstrates a low response, and Subjects D2-2, D2-6, and D2-4 illustrate medium responses that correspond to relatively high probability values for all involved measurements. We speculate that Subject D2-1 had a missed second infection because their antibody response defies the trend, increasing with days post infection. In Fig. 19(b), Subject D2-9 exhibits a high response, Subject D2-7 demonstrates a medium response, and Subject D2-9 illustrates a low response. Subject D2-8 shows a response that drops sharply, defying the trend of slow decay. While we do not expect missed events with vaccination, since inoculation doses are well-recorded, there is population variability in the durability of antibody response to a vaccine.
Figure 20: Log-transformed antibody measurements from the infected, then vaccinated (VI) population from Dataset 2 with personal trajectories of interest with corresponding probability models. Measurements with recorded monotonically decreasing OD values have been AUC-extrapolated using the next titration level. See Supplementary Material for accompanying video.
@@
For vaccination after infection, in Fig. 20, we see that Subject D2-11 exhibits a medium response to both infection and vaccination, whereas a low response to both is observed for Subject D2-14. Subjects D2-12 and D2-15 have above-average responses to both infection and vaccination, and in particular, Subject D2-12 exhibits an increasing trajectory, which is against the trend and prediction of our model. Finally, Subject D2-13 shows a low response to infection, but responds well to vaccination.
@@ 8.5 Comparison of Datasets 1 and 2
Datasets 1 and 2 were recorded using different measurement systems [22, 23, 24]; the first was measured in IU/mL, and the second in unitless AUC. Even after our titration-extrapolation for Dataset 2 (Section 6.2), the antibody decay is more pronounced or visible in Dataset 1 than Dataset 2, as evidenced by the $k_{1}$ and $k_{2}$ values shown in Tables 1 and 2; this is likely due to the varied measurement recording systems. These $k_{1}$ and $k_{2}$ parameters are proxies for decay speed. For Dataset 1, we find a faster decay for infection than vaccination, but the reverse is true for Dataset 2. We find the vaccinated measurements to be more clustered across time for Dataset 2, as evidenced by the fact that 97 % of measurements fall within the model contours, with 67 % at a “high” probability level, as compared to 71 % and 49 % for Dataset 1. For infection: 60 % are at a “high” probability level for Dataset 1, whereas only 21 % of measurements are for Dataset 2, indicating greater variability in the second population’s response to infection.
@@ 8.6 Limitations and extensions
A remaining assumption in our modeling is that the entire population starts in the naïve state, which is appropriate for characterizing an emergent disease. This simplifies some of our work, including the ability to set the disease and vaccination prevalence on day zero to zero. In the future, one could design a framework for a currently circulating disease, which could be validated using real incidence rates and antibody measurements.
In future work, we will employ the absolute and the personal timeline models to represent the relevant probabilities for the general, time-inhomogeneous setting by enumerating all possible personal trajectories, which will likely involve a challenging application of enumerative combinatorics. Some guiding principles will inform this process. Due to the underlying biology, extreme sequences of events, such as an individual becoming newly infected every day, should be assigned very low probabilities. Further, despite the multitude of potential sequences of events leading to an antibody response and current state on a particular time step, the likelihood of infection or vaccination on the next time step depends solely on the current state information.

### v2

@@ 7 Example applied to population transmission
While the previous section was focused on antibody responses in individuals, the transmission of a disease through a population is another important factor that governs the total antibody distribution in a population over time. We now turn our attention to the different possible personal event sequences for individuals as an infection and its vaccination transmits through a population. We consider two separate time-homogeneous transition matrices and analyze their behavior over 505050 time steps, or around 222 years.
@@
We first consider $S_{1}$ (Figure 8a) representing a disease spread with a rate of infection that is a magnitude higher than that of vaccination. Recall that our states are ordered as $N,\mathcal{I},\mathcal{V},\mathcal{I}^{\prime},\mathcal{V}^{\prime}$ in all transition probability matrices with the column representing the transition out of the corresponding state. We observe (Figure 8a) that due to larger probabilities of infection, the median number of infections per individual is 444 and the median previously infected status is maintained for 353535 time steps. In contrast, the vaccinations have medians of 0 and naïve status is maintained for a median of 888 time steps. Some representative individual event sequences are demonstrated in Figure 8b, each row represents a personal trajectory for an individual, the colors represent the immune status for the individual at the time step corresponding to the column.
@@
Figure 8: Simulations for a population with 100010001000 individuals with a time-homogeneous transition matrix $S_{1}$, per 505050 time steps, i.e., approximately a two year period. (a) Histograms of counts of total naïve, infections, vaccinations, previously infected status, and previously vaccinated status. (b) Three representative sequences. Light yellow denotes the naïve state (1, $N$), red and pink the newly and previously infected states (2, $\mathcal{I}$ and 4, $\mathcal{I}^{\prime}$), and blue and light blue the newly and previously vaccinated states (3, $\mathcal{V}$ and 5, $\mathcal{V}^{\prime}$).
@@
We increase the probability of transition to naïve from both previously infected and previously vaccinated status to obtain $S_{2}$ (Figure 9a). This emulates diseases with shorter-term immunity and thus a quicker return to the naïve status; see Figure 9a. As anticipated, this increases the likelihood of sequences with a higher number of immune events; see Figure 9b.
@@
Figure 9: Simulations for a population with 100010001000 individuals with a time-homogeneous transition matrix $S_{2}$, per 505050 time steps, i.e., approximately two year period. (a) Histograms of counts of total infections, vaccinations, previously infected status, previously vaccinated status, and naïve status. (b) Five representative sequences. Light yellow denotes the naïve state (1, $N$), red and pink the newly and previously infected states (2, $\mathcal{I}$ and 4, $\mathcal{I}^{\prime}$), and blue and light blue the newly and previously vaccinated states (3, $\mathcal{V}$ and 5, $\mathcal{V}^{\prime}$).
@@ 8 Discussion
In this section, we discuss a generalization of the two-event time-inhomogeneous model, more on the general time-inhomogeneous model, and then investigate the variability of antibody response across a population. We include a comparison of the two datasets used in our validation of our methods, and finish with limitations and extensions of our work, as well as implications for immunologists.
@@ 8.2 General time-inhomogeneous model
Towards constructing the conditional probability models in the general setting of the time-inhomogeneous model, we discuss an important check. We want to ensure that the number of high-probability sequences with a large number of events is vanishingly small to reflect real world situations, as a person will not realistically become reinfected or revaccinated at every point in time. Notice that with the possibility of either infection or vaccination at every time step, we obtain $2^{T}$ possible event sequences over $T$ time steps. In reality, we can constrain this number by quite a bit. We note that as class $\mathcal{I}$ necessarily transitions to class $\mathcal{I}^{\prime}$ and class $\mathcal{V}$ to class $\mathcal{V}^{\prime}$, there are at the most $\lfloor T/2\rfloor$ events possible before time $T$, which means there are approximately $2^{\lfloor T/2\rfloor}$ ways for events to occur before time $T$. We recall also that time will likely be batched by several weeks, so $T=10$ might mean 210 days if $\Delta T=21$ days or 3 weeks, which is a reasonable time spacing. Then for this example, there would be $2^{5}=32$ possible sequences in 210 days, which is a manageable number. We can further constrain this set by calculating the probability of obtaining such a sequence as a product of the associated transition probabilities. What we need to additionally consider are the days spent between two immunological events in the previously infected or previously vaccinated state as this affects the antibody response distribution. As an illustration, consider $T=6$. For simplicity, let us assume that $z_{i}\in\{\mathcal{I},\mathcal{V}\}$. The 43 total trajectories are:
@@ 8.3 Personal trajectories of interest
We study personal trajectories of interest for a few subjects from both datasets. To discuss individual subjects, we standardize and anonymize the IDs by dataset (D1 or D2 for Dataset 1 or Dataset 2) and appearance number in the following figures by dataset (e.g., D1-1). In Figure 10, we plot the single-event models for infection and vaccination with representative and outlier personal trajectories of interest.
@@
Figure 10: Log-transformed antibody measurements from the infected (I) (a),(b) and vaccinated (V) (c),(d) populations from Dataset 1 (a),(c) and Dataset 2 (b),(d) with personal trajectories of interest with corresponding probability models. For Dataset 2, measurements with recorded monotonically decreasing OD values have been AUC-extrapolated using the next titration level.
Single infection.
For Dataset 1, in Figure 10a, we see that Subject D1-5 maintains a high response and Subjects D1-1, D1-3, and D1-4 illustrate medium responses after the initial acute phase that correspond to relatively high probability values for all involved measurements. However, Subject D1-1 shows a rapid but medium response initial increase, Subject D1-4 has a rapid, large initial increase, and Subject D1-3 has a slower and medium response initial increase in antibody level. Subject D1-2 has a rapid, high initial response and then a quick decay to a below average antibody level.
@@
For Dataset 2, in Figure 10b, Subject D2-5 exhibits a high response, Subject D2-3 demonstrates a low response, and Subjects D2-2, D2-6, and D2-4 illustrate medium responses that correspond to relatively high probability values for all involved measurements. We speculate that Subject D2-1 had a missed second infection because their antibody response defies the trend, increasing with days post infection.
Single vaccination.
Missed events are unexpected for vaccination since immunization doses are well-recorded, but there is still population variability in the durability of antibody response to a vaccine. For Dataset 1, in Figure 10c, we see that Subject D1-8 shows a medium response after the initial vaccination. Subjects D1-10 and D1-11 are low responders, as Subject D1-11 never mounts much of an antibody response at all, and Subject D1-10’s response falls below the lowest model contour before 200 days after vaccination. Subjects D1-7 and D1-9 have relatively high initial responses, but Subject D1-7’s antibody level falls much more quickly than Subject D1-9. Subject D1-6 exhibits a medium initial response, but this is sustained better than many subjects beyond 100 days, and then we speculate that they experienced a missed infection sometime before their next measurement (over 400 days after vaccination).
For Dataset 2, in Figure 10d, Subject D2-9 exhibits a high response, Subject D2-7 demonstrates a medium response, and Subject D2-9 illustrates a low response. Subject D2-8 shows a response that drops sharply, defying the trend of slow decay.
Two vaccinations.
In Figure 11 we plot individualized two-event models for a booster dose after vaccination for six subjects of interest from Dataset 1. Subject D1-15 exhibits a medium response to both infection and vaccination, with representative decay patterns that are well captured by our model. This is true for Subject D1-16 for the measurements up until the booster dose, roughly 100 days after which we speculate that a breakthrough infection occurred, since the subject’s antibody response spikes again. We observe that Subject D1-17 mounted a relatively low response to the first vaccine sequence, but a stronger response was induced by the booster dose. Subject D1-18 exhibits a low initial vaccine response that drops quickly, followed by a medium response to the booster that also rapidly drops. Finally, Subjects D1-20 and D1-19 show low and high antibody responses to both the initial and booster vaccinations, respectively, with a particularly high and sustained response to the booster dose for Subject D1-19.
Figure 11: Log-transformed antibody measurements from the vaccinated then boosted (VV) population from Dataset 1 with personal trajectories of interest with corresponding probability models. The relative date of the booster dose is marked with a vertical, white dashed line. See this Supplementary Figure (hyperlink) for the corresponding video and Appendix E for a description.
Vaccination after infection.
@@
For Dataset 2, in Figure 12, we see that Subject D2-11 exhibits a medium response to both infection and vaccination, whereas a low response to both is observed for Subject D2-14. Subjects D2-12 and D2-15 have above-average responses to both infection and vaccination, and in particular, Subject D2-12 exhibits an increasing trajectory, which is against the trend and prediction of our model. Finally, Subject D2-13 shows a low response to infection, but responds well to vaccination. A similar analysis for subjects from Dataset 1 is shown and discussed in Appendix C (see Figure C1).
Figure 12: Log-transformed antibody measurements from the infected, then vaccinated (VI) population from Dataset 2 with personal trajectories of interest with corresponding probability models. Measurements with recorded monotonically decreasing OD values have been AUC-extrapolated using the next titration level. See this Supplementary Figure (hyperlink) for the corresponding video and Appendix E for a description.
@@ 8.4 Comparison of Datasets 1 and 2
Datasets 1 and 2 were recorded using different measurement systems [22, 23, 24]; the first in IU/mL, and the second in unitless AUC. Even after titration-extrapolation for Dataset 2 (Section 6.2), the antibody decay is more pronounced or visible in Dataset 1 than Dataset 2, as evidenced by the $k_{1}$ and $k_{2}$ values shown in Table 1 and Table 3 ($k_{1}$ and $k_{2}$ are proxies for decay speed); this is likely due to the varied measurement recording systems. For Dataset 1, we find a faster decay for infection than vaccination, but the reverse is true for Dataset 2. We find the vaccinated measurements to be more clustered across time for Dataset 2, as evidenced by the fact that 97 % of measurements fall within the model contours, with 67 % at a “high” probability level, as compared to 71 % and 49 % for Dataset 1. For infection: 60 % are at a “high” probability level for Dataset 1, whereas only 21 % of measurements are for Dataset 2, indicating greater variability in the second population’s response to infection.
@@ 8.5 Limitations and extensions
A remaining assumption in our modeling is that the entire population starts in the naïve state, which characterizes an emergent disease. This simplifies some of our work, including the ability to set the disease and vaccination prevalence on day zero to zero. In the future, one could design a framework for a currently circulating disease, which could be validated using real incidence rates and antibody measurements.
In future work, we will employ the absolute and the personal timeline models to represent the relevant probabilities for the general, time-inhomogeneous setting by enumerating all possible personal trajectories, which will likely involve a challenging application of enumerative combinatorics. Some guiding principles inform this process. Due to the biology, extreme sequences of events, such as an individual becoming newly infected every day, will be assigned very low probabilities. Further, despite the multitude of potential sequences of events leading to an antibody response and current state on a particular time step, the likelihood of infection or vaccination on the next time step depends solely on the current state information.